OpenStudy (anonymous):

f(x)=x^2+4 and g(x)=x+5 find the following (gof)(x)

OpenStudy (amistre64):

place f(x) INTO g(x) (f(x))+5

OpenStudy (anonymous):

To solve this problem, you just need to do some simple substitution. The notation (gof)(x) means finding a new expression for g(x) when f(x) is plugged in as the argument (instead of just "x"). So \[(fog)(x)=(x+5)^2+4 \] Note that we've plugged g(x)=x+5 in for x in the expression for f(x)! And \[(gof)(x)=(x^2+4)+5=x^2+9\] Here, we've plugged in f(x) for x in the expression for g(x). I hope that solidifies your understanding!

OpenStudy (anonymous):

That helps alot, I see how you did that thanks